OpenStudy (anonymous):

HELP PLZ! I DONT GET THIS The area of a rectangular piece of land is 280 square meters. If the length of the land was 5 meters less and the width was 1 meter more, the shape of the land would be a square. Part A: Write an equation to find the width (x) of the land. Show the steps. Part B: What is the width of the land in meters? Show steps.

OpenStudy (anonymous):

if we call length=a and width=b then axb=280 then i can say that a-5=b+1 because it is a square now. a=b+6 then (b+6) x b = 280 b^2 + 6b = 280 b=14

OpenStudy (anonymous):

Thanks, so for part A i put this: W = width of rectangle L = length of rectangle L*W = 280 So we know that if we where to change the length to be 5 meters less therefore the width to be 1 meter more, they will be equal. L − 5 = w + 1 but then I'd change l and w to a and b?

OpenStudy (anonymous):

Then part B just show the work right?

OpenStudy (anonymous):

And where did you get 6?

OpenStudy (anonymous):

L-5 = W+1 so we can say L=W+6 L*W = 280 so we can change L and w+6 because they are equal. (W+6)*W=280 W=14

OpenStudy (anonymous):

Thank you!!